Yaamini’s Notebook: Manchester Conditioning Update

It’s been a slow burn

As I expected, there were problems getting the heater to warm the water in the header tank fast enough to meet the desired temperature schedule I set. Laura mentioned that the temperature fluxuated during the day, so she was wary of increasing the temperature.

Table 1. Actual water temperatures read from AVTECH. Temperature listed was the peak temperature that day.

Day Date Temperature
6 6/16/17 17.06 ºC
7 6/17/17 17.31 ºC
8 6/18/17 18.56 ºC
9 6/19/17 18.50 ºC

The temperature is holding around 18.5 ºC, so I increased the temperature setting on the AVTECH two degrees. My goal is to get the temperature up to 20ºC. This will only set back my conditioning timeline two days!

Table 2. Desired temperatures for the remainder of conditioning.

Day Date Temperature
10 6/20/17 20 ºC
11 6/21/17 21 ºC
12 6/22/17 22 ºC
13 6/23/17 23 ºC
14-26 6/24/17-7/6/17 23 ºC

I’ll be back on Day 13 to check on things and Laura will continue to raise the set point and check the AVTECH.

from the responsible grad student http://ift.tt/2rNUlUC
via IFTTT

Sean’s Notebook: Trimming and Quality…

Sean’s Notebook: Trimming and Quality Checking EPI-135 and EPI-135WG.

We’re back to playing with Hollie’s Geoduck methylation data, and noticed that there was some wonky nucleotide sequence at the beginning of reads potentially hampering the mapping efficiency downstream. Steven asked me to trim the first 18 nucleotides and last nucleotide from each sequence, and I did that here.

EPI-135 Pre Trimming:
Read 1:
EPI-135_1-untrimmed

Read 2:
EPI-135_2-untrimmed

EPI-135 Post Trimming
Read 1:
EPI-135_1-Untrimmed

Read 2:
EPI-135_2-Untrimmed

EPI-135WG Pre Trimming:
Read 1:
EPI-135WG_1-UnTrimmed

Read2:
EPI-135WG_2-UnTrimmed

EPI-135WG Post Trimming:
Read 1:
EPI-135WG_1-Trimmed

Read 2:
EPI-135WG_2-Trimmed

It looks like trimming cleaned most everything up, though maybe it could have stood to have an additional base pair or two trimmed off the tail.

I’ve started running Bismark on the outputs to see if this trimming improves mapping rates, hopefully it will. I’m going to try tinkering with the number of allowed mismatches in Bowtie, to see if that helps also.

On the Oly Genome front. I talked with Katherine today, and have a pretty decent idea of a game plan for going forward. It looks like I was under the wrong impression that all the Illumina data was paired end, it looks like the longest insert stuff was actually mate pair, which the Platanus devs say is not ideal for assembling. Oops.

The plan:

1. Finish polishing the Canu assembly with Pilon. Still mapping reads back to the Canu assembly with bowtie.
2. Re-run Platanus using only the PE150 data from Illumina, then scaffold with PE150 and MP50 data. Then throw it in to Redundans **without** the reduction step. Katherine thinks that Redundans may be throwing away too much data due to high heterozygosity, and turning that step off may prevent the loss.
3. Throw the BGI assembly, the Platanus/Redundans Assembly, and the Canu/Pilon assembly in to a meta assembler such as GARM.
4. Have the best assembly ever. Or at least a better assembly.

Grace’s Notebook: June 19, 2017 Manchester

Helped Laura today at Manchester along with Olivia.

I screened the larvae at 224, 180, and 100. Laura counted the larvae. Olivia brought the buckets over, switched out droppers and air stones, and cleaned.

After all that, I screened the new larvae from the oysters that spawned on 100 micron screens. These were placed in tri-pours and counted. Then, it was determined how much of the water containing the larvae should be re-stocked in the large buckets in order to stock 50K larvae. The remaining larvae were preserved in ethanol.

Laura’s Notebook: Manchester, 6/19

Screening day and more! Things went very smoothly today. Grace, Olivia and I make a good team.

Arrival inspection:

  • Banjos very dirt, a few pseudo-clogged. Due to heavy use of Tet algae.
  • HL-10 Ambient, which spawned, banjo overflowing but luckily lots was caught in an empty trio pour that was sitting adjacent to the catchment bucket

Screened larvae (Olivia, Grace, and me).

  • Mortality ranged dramatically between groups. Very high mortality in the following groups: – SN-6 Low 180um – lots of ciliates present – SN-10 Low 224um, 180um – lots of ciliates present – NF-10 Low 224um, 180um

imageimage

Collected new larvae, counted, stocked, and sampled. Grace did the majority of this process today.

6d0a3b25-93a2-4c00-ada2-d626b2cc58ca

Cleaned gigas (Olivia)

  • Drained all three tanks, cleaned with vortex- LOTS of stringy stuff in the top tank w/o animals. I now am certain that the mucus is a result of the Reed’s paste, and not coming from the animals: file_009
  • Rinsed gigas, checked for morts – there were none.
  • Refilled tanks, then turned off flow and allowed header temp to increase.
  • Grace adjusted flow for 1.2 L/Min on each side.
  • Increased set point temp to 21degC.
  • Fed with 490mL Reed’s paste
  • Installed new pump (had been using small PSRF pump).

Cleaned Broodstock (Olivia):

  • First, recorded broodstock location on manifold
  • [image here]
  • Dumped all broodstock & larval catchment buckets, rinsed oysters, cleaned buckets with vortex.

Larval experiment water change:

Katherine recommends having 5 replicates.

  • [insert Katherine’s post here]
  • I heeded her advice, and thus split my 3 reps into 5 via the following:
    • Math!:
      • Assuming 800 larvae in each tripour, total # larvae per treatment = 800*3 = 2,400
      • Larvae per rep for 5-reps = 2,400/5 = 480
      • larvae to remove from each tripour = 800-480 = 320. Split these up into 2 new silos = 160 larvae per new silo per triplicate
    • How this was done:
      • Set 2 new silos in FSW. Label with designated treatment group.
      • Working 1 silo at a time, transfer larvae to a tripour, fill to 200mL
        • mL needed per new silo = 800 larvae/200 mL = 4 larvae/mL, 160 larvae / l/mL = 40 mL
      • Plunge 20 times, pour 40mL into 2 falcon tubes. Add the contents into the two silos.
      • Repeat with other 2 triplicates.
      • Result:
        • 2 new silos will have 160*3 larvae = 480
        • Original silos will have 800-(2*160) = 480
    • New setup: file_000 4
    • Changed water.
      • Filled a 5-gal bucket with 18L.
      • Harvested 450mL Tiso, CGW & counted density:

1ca27f50-b20e-460c-9d61-ceb11e763dbc

  • Mimicking Katherine’s experiment, I will sample & image larvae on Thursday 6/22 (day 7), then also on 6/29 (day 14)
  • Reminder: I’m working with the 4 treatment groups in the South Sound (SN) population.

current summary data:

imageimageimageimage

from The Shell Game http://ift.tt/2tIbLyt
via IFTTT

Laura’s Notebook: Manchester, 6/18

Today’s vista: file_007file_008

Water change on larval experiment

  • Harvested 450 mL Tiso + 450 mL Chagra
  • Counted algae. Here’s what they look like under the scope. I count 5 out of the 9 large squares in the hemocytometer grid (each large square contains 16 small squares). This first photo shows one large square, scope is at 10x: file_000

And here’s algae at a higher lens; I believe this is 30x, but will need to confirm: file_002

Counts: a44fec13-6ad3-4789-a768-781e5224e140a44fec13-6ad3-4789-a768-781e5224e140

  • Filled 5gal bucket with 13L FSW, added 690 mL algae cocktail, plunged and distributed among 12 clean tripours
  • Transferred larval silos to new medium

Changed banjos in SN & NF groups (hadn’t done that yesterday)

Imaged all larvae screened on 6/15. These were not killed with lugols, and were kept in the fridge. Changed up my imaging protocol: – Used the dissecting scope. I think the photos are much nicer: file_006file_005

 - Imaged a ruler (mm) at 12x and 32x ![file_003](http://ift.tt/2rykJ0h) ![file_004](http://ift.tt/2sRC3kS) - Imaged all wells at at 32x. - New protocol for imaging will be: - Don’t add lugols - Store in fridge - image 1-3 days post collection - Use dissecting scope, image only at 32x, image as many larvae as possible. Larvae collect in the center of the well, in a cluster, so should be able to capture most in one image. Fed - Setting/K&HL broodstock table: - 1/2 609 - 1/4 Tet - 1/4 Chagra - Larval table/SN&NF broodstock table: - 1/2 Tet - 1/2 Chagra Gigas: - Increased set point temp to 21degC. - Since temp is not achieving set point, I tried switching to the heated line, but the flow on the heated line was not sufficient enough to achieve 2.4L/min. Swapped back to the ambient line, and adjusted flow to achieve 2.4L/Min. - **When swapping lines on gigas, the high pressure on the ambient line blew the tube off the labcock valve, which sent a ton of ambient, very dirty water through and into one of my 100um larval buckets - NF-6 Ambient. I likely lost some larvae, and the bucket was extremely dirty. I screened this bucket through 200 onto 100um, trying to remove some gunk. I did, but not all of it. I returned the remnants back to the larval bucket. Luckily, the flow had been pushing live larvae into the 180um bucket since yesterday afternoon, and there were very few 100um larvae in this bucket to start with, only ~2,700: - Screening counts: ![2f62fbc7-99c4-4969-9e55-59741ad20f3d](http://ift.tt/2ryoRxk) - And none stocked since 6/15:  

8e39bf5b-9ea9-460b-ac1f-634763f160cb

from The Shell Game http://ift.tt/2ry5zZd
via IFTTT

Laura’s Notebook: Manchester, 6/16 & 6/17

6/16

  • Counted rest of larvae from 6/15 screening
  • Imaged all new larvae from 6/12, 6/14 & 6/15
  • Prepared FSW+algae for larval experiment water change:
    • Filled Harvested 400mL Tiso & 400mL Chagra. I want 100k cells algae/mL in my larval experiment tripours, which are filled to 800mL. To determine algae concentration, I plunged and sampled 20ul and injected into hemocytometer.
    • Counted 5 of the 1mm^2 squares. Here are my calcs from the past few days:

6/17

Arrived, collected new larvae, of which there were lots:

image

Noticed that the algae dosing pump @ the small table (setting tank/K & HL broodstock) was not turned on. They were without food since X, so for ~X hrs. I fed each setting silo 200mL of the algae cocktail in the algae header.

Prepared FSW+algae for larval experiment water change:

  • I used a new water pre system, based on Katherine’s protocol from last summer:
  • Filled a bucket with 13L FSW (taken from my small table manifold, with the algae pump turned off).
  • Harvested 400mL Tiso & 400mL Chagra. I want 100k cells algae/mL in my larval experiment tripours, which are filled to 800mL. To determine algae concentration, I plunged and sampled 20ul and injected into hemocytometer.
  • Counted 5 of the 1mm^2 squares. Here are my calcs from the past few days: image
  • Mixed 630mL algae cocktail into 13L FSW, plunged, and filled 12 tripours to 800mL.
  • Transferred larval silos into fresh water.

Larval bucket maintenance:

  • Set up 2-bucket challenge by connecting the 100um and 180um within the SN & NF groups, and added a second bucket for the HL & K groups. This ends up taking ~1-1.5hrs; should determine how I can streamline this task.

Cleaned outside upwelling tank: turned of inflow, drained, filled with fresh water & added 400mL Vortex. Left pump on. Will leave it overnight to clean.

Gigas maintenance:

  • Vacuumed as best I could. Noticed that there was lots of mucus in the east tank.
  • Moved temperature probe & Durafet 3 into the east tank.
  • Increased immersion heater set point to 20degC..
  • Filled algae header with 485mL Reeds + fresh water.

from The Shell Game http://ift.tt/2rL40uR
via IFTTT

Sean’s Notebook: Canu run finished

Looks like Hyak finished it’s Canu assembly for the Oly PacBio data over the weekend, and here are the statistics for it.

[seanb80@mox2 oly-test]$ /gscratch/srlab/programs/canu/Linux-amd64/bin/tgStoreDump -sizes -s 50000000 -T /gscratch/srlab/data/CanuTest/oly-test/unitigging/oly_pacbio_.ctgStore 2 -G /gscratch/srlab/data/CanuTest/oly-test/unitigging/oly_pacbio_.gkpStore
lenContig ng10       23155 bp   lg10     175   sum    5001399 bp
lenContig ng20       19201 bp   lg20     416   sum   10012857 bp
lenContig ng30       16610 bp   lg30     697   sum   15014659 bp
lenContig ng40       15073 bp   lg40    1014   sum   20013348 bp
lenContig ng50       13606 bp   lg50    1364   sum   25008188 bp
lenContig ng60       12421 bp   lg60    1749   sum   30009994 bp
lenContig ng70       11361 bp   lg70    2170   sum   35007399 bp
lenContig ng80       10115 bp   lg80    2634   sum   40002506 bp
lenContig ng90        8041 bp   lg90    3178   sum   45002283 bp
lenContig sum   46288879 (genomeSize 50000000)
lenContig num       3388
lenContig ave      13662

13.6 kbp NG50 is an improvement! Next, polishing with Pilon.